# On Selection of Cross-Section Averages in Non-stationary Environments

### v3

@@ Abstract
Information criteria (ICs) have been widely used in factor models to estimate an unknown number of latent factors. It has recently been shown that ICs perform well in Common Correlated Effects (CCE) and related settings when selecting a set of cross-section averages (CAs) sufficient for the factor space under stationary factors. As CAs can proxy non-stationary factors, it is tempting to claim an excellent performance of ICs under general factors, too. We show formally and in simulations that they remain consistent, but the more persistent factors are, the poorer they perform in small samples, which goes against the sentiment in the CCE literature.
@@ 1 Introduction
where $\mathbf{Z}_{i}=[\mathbf{z}_{i,1},\ldots,\mathbf{z}_{i,T}]^{\prime}\in\mathbb{R}^{T\times K}$, $\mathbf{U}_{i}$ is an error term, while $\mathbf{F}\in\mathbb{R}^{T\times m}$ are latent common factors, where $m$ denotes the number of factors, and $\mathbf{C}_{i}\in\mathbb{R}^{m\times K}$ is the loading matrix. The most popular setting nested in (1.1) is that of interactive effects. Then $\mathbf{Z}_{i}=[\mathbf{y}_{i},\mathbf{X}_{i}]$, and $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ gives a model with unobserved heterogeneity, where $\mathbf{X}_{i}$ is a $T\times k$ matrix of explanatory variables. Naturally, $K=k+1$, and $i$ indexes individuals, such as countries or households. The unobserved component of $\mathbf{y}_{i}$ is likely to be correlated with $\mathbf{X}_{i}$ (see Bai, 2009a, ). If $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$, (1.1) represents the Common Correlated Effects (CCE) setting of Pesaran, (2006), where $\boldsymbol{\beta}\in\mathbb{R}^{k}$ is the parameter of interest, $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i}$, $\boldsymbol{\gamma}_{i}$) are the idiosyncratics (loadings). Since both $\mathbf{X}_{i}$ and $\mathbf{y}_{i}$ load on $\mathbf{F}$, we have endogeneity and cross-section dependence among individuals that make the least squares (LS) estimator of $\boldsymbol{\beta}$ inconsistent.
@@
It is possible to estimate unobserved factors in several ways, including the Principal Components (PC) method (see Bai and Ng,, 2006) or diversified projections (see Fan and Liao,, 2022). Pesaran, (2006) suggests a simple and elegant way to proxy $\mathbf{F}$ up to a linear transformation by taking cross-section averages (CAs): $\overline{\mathbf{Z}}=[\overline{\mathbf{y}},\overline{\mathbf{X}}]=\widehat{\mathbf{F}}=\mathbf{F}\overline{\mathbf{C}}+o_{p}(1)$, because CAs of the idiosyncratics are negligible under a wide range of empirically relevant assumptions (see e.g. Pesaran and Tosetti,, 2011). The CCE estimator is then given by
$\widehat{\boldsymbol{\beta}}_{CCE}=\overline{\mathbf{Q}}^{-1}\overline{\mathbf{q}},$ (1.2)
where $\overline{\mathbf{Q}}=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}}\mathbf{X}_{i}$, $\overline{\mathbf{q}}=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}}\mathbf{y}_{i}$, $\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix and $\mathbf{A}^{+}$ is the Moore-Penrose inverse. The core principle of CCE is to project the variables onto the orthogonal complement of their CAs in order to purge the factors asymptotically. Note that CAs are applicable beyond the usual interactive effects and CCE. Stauskas and De Vos, (2025) consider Distinct Correlated Effects (DCE), where $\mathbf{y}_{i}$ and $\mathbf{X}_{i}$ load on distinct sets of factors, and (1.2) is aided by bootstrap. Massacci and Kapetanios, (2024) use $\overline{\mathbf{X}}$ to approximate factors and produce factor-augmented forecasts.
@@
A substantial advantage of CAs is their ability to estimate $\mathbf{F}$ irrespective of their time series properties. Indeed, the CCE estimator enjoys popularity, because it is consistent and asymptotically normal under very general $\mathbf{F}$ in large $N,T$ settings without any modifications (see e.g. Kapetanios et al.,, 2011, Westerlund,, 2018, Westerlund and Petrova,, 2018, or Stauskas,, 2023). This property is valid if the rank of $\overline{\mathbf{C}}$ is equal to $m$, which means that the number of factors cannot exceed the effective number of CAs. However, there are reasons to ensure that $m=k+1$. If $m<k+1$ and $TN^{-1}$ is bounded, CCE produces an asymptotic bias whose analytical correction is infeasible (see Karabiyik et al.,, 2017). This situation is common in macro datasets (see De Vos and Stauskas,, 2024). In general, too many CAs might even lead to a model that over-fits the data. In contrast, inclusion of too few generally leads to inconsistent estimates of the model parameters (see Juodis,, 2022, for a discussion). Beyond CCE, Stauskas and De Vos, (2025) show that DCE only uses $\overline{\mathbf{X}}$, but $m=k$ must hold. In forecasting, the same is needed to avoid only conservative confidence intervals around factor-augmented forecasts (see Karabiyik and Westerlund,, 2021). In response to this, Margaritella and Westerlund, (2023) (MW) were the first to demonstrate that the information criteria (ICs) of Bai and Ng, (2002) and Bai, 2009b from the PC literature can be applied in the CCE setting under stationary factors to identify an optimal set of CAs from $\overline{\mathbf{Z}}$. De Vos and Stauskas, (2024) (DVS) show similar results when CAs are selected from $\overline{\mathbf{X}}$, which is more applicable in the DCE and forecasting settings. [^1] Even in basic CCE setting, it is advised to omit $\overline{\mathbf{y}}$ and use $\overline{\mathbf{X}}$ only to avoid computational issues (see Karavias et al.,, 2023)., [^2] PC techniques are limited in CCE, as they focus solely on $m$ by detecting the largest eigenvalues of the data matrix. They cannot detect sets of CAs as they do not have a natural ordering. With IC, we learn $m$ from the cardinality of the selected set.
@@
MW stress that an assumption of stationary $\mathbf{F}$ is required only to simplify the proofs, hinting at a much greater generality of IC. As CAs proxy a general factor structure, it is natural to evaluate this statement and re-visit the ICs proposed in MW and DVS. Therefore, this study can be seen as the CAs counterpart of Bai, (2004), where classical PC results of Bai and Ng, (2002) are evaluated against pure unit root factors. Instead, we use a mildly integrated process by Magdalinos and Phillips, (2009) to model $\mathbf{F}$, which allows us to experiment with varying degrees of factor persistence. We demonstrate formally and in simulations that, while ICs remain consistent as $(N,T)\to\infty$, highly persistent factors negatively affect their small sample performance. We discuss differences between our results and those in Bai, (2004), and also explore penalty functions adapted to non-stationary $\mathbf{F}$, as suggested in the latter study. Since they render our ICs inconsistent, we argue that practitioners should not mechanically take recommendations from the PC literature if the goal is to select CAs. More importantly, the ICs of DVS and MW (with the original penalties) should be applied in relatively large samples if the presence of non-stationary factors is suspected.
@@ 2 Econometric Setup
We focus on the IC of DVS due to its wider applicability. While we provide numerical experiments on both DVS and MW, theoretical arguments regarding the latter are nearly identical, and we relegate them to Section 3.3 of the Supplementary material. To operationalize our analysis, we introduce $M$ which is a set of column indices of $\overline{\mathbf{X}}$, and $\mathbf{q}_{M}\in\mathbb{R}^{k\times g}$ picks the corresponding $g$ averages in practice. For instance, $\overline{\mathbf{X}}\mathbf{q}_{M}=\widehat{\mathbf{F}}_{M}$ selects $g$ out of $k$ CAs. Consequently, let $M_{0}$ denote the true set of averages from $\overline{\mathbf{X}}$ such that $\mathrm{rank}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}})=m$, and $|M_{0}|=m$, where $|M|$ denotes the cardinality of an arbitrary set $M$. IC under consideration is given by
@@
where $\ln(.)$ denotes the natural logarithm and $\mathrm{det}(\mathbf{A})$ is a determinant of any square matrix $\mathbf{A}$. $\overline{\mathbf{Q}}_{M}$ is $\overline{\mathbf{Q}}$ from (1.2) evaluated at $\widehat{\mathbf{F}}_{M}$, and $p_{N,T}$ is a penalty term. [^3] In case of MW, $\overline{\mathbf{Q}}_{M}=\ln\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M}}\widehat{\boldsymbol{\nu}}_{i}\right)$, where $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ and $\widehat{\boldsymbol{\beta}}$ is obtained using the CCE estimator under all available $k+1$ CAs. Examples of feasible and most popular penalties are given by
@@
for $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$ (see more in Section 5 of Bai and Ng,, 2002). Note that (2.1) is function of (a version of) the denominator of the CCE estimator, which is robust to general unknown factors (see Theorem 1 in Westerlund,, 2018).
@@
where $|\widehat{M}|=g$ provides the estimator of the number of factors as a consequence. In DVS, it is demonstrated that $\mathbb{P}(\widehat{M}=M_{0})\to 1$ as $(N,T)\to\infty$. This result theoretically justifies the use of IC previously very popular in the PC literature to select the number of factors. These findings are based on stationary $\mathbf{F}$, which has been considerably relaxed in the CCE/CAs literature. Therefore, our natural goal is to examine (2.1) under departures from this restrictive assumption. Throughout our analysis, we employ the following set of assumptions.
Assumption 1. $\{\mathbf{f}_{t}\}$ is a mildly integrated process as defined in Magdalinos and Phillips, (2009).
  • (a)
    In particular,
    $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{G}T^{-\tau},\quad\mathbf{G}>0\hskip 5.69054pt(diagonal),$
@@
  • (b)
    If $\tau=0$, then $\{\mathbf{f}_{t}\}$ is a stationary process with summable autocovariances.
@@
Assumption 1 (a) treats the factors very flexibly and offers comparative statics, as $\tau\to 1$ increases persistence of the process. We are not interested in a specific model for $\mathbf{F}$, but Magdalinos and Phillips, (2009) allow us to vary $\tau$ and give theoretical guarantees, such as $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{F}\to_{p}\boldsymbol{\Sigma}_{\mathbf{F}}$, which is a constant positive definite matrix. Part (b) ensures that we return to the usual stationarity conditions under $\tau=0$. Specifically, we need the eigenvalues of $\mathbf{I}_{m}-\mathbf{G}$ to be within a unit circle, so that the process can be inverted and admit a representation of $MA(\infty)$. Assumption 2 (a) is split into two parts. If the idiosyncratics $\{\mathbf{e}_{i,t}\}$ are correlated over time, for $\mathbf{E}_{i}=[\mathbf{e}_{i,1},\ldots,\mathbf{e}_{i,T}]^{\prime}$, we have that $\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}\right\|=O_{p}(T^{-\tau/2})$ (see Lemma 3.1 in Magdalinos and Phillips,, 2009), which is too slow to demonstrate that the factor estimation error is negligible. In order not to obscure the very effect of factor persistence on IC, we restrict the correlation of idiosyncratics and show that $\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}\right\|=O_{p}(T^{-(1+\tau)/2})$ as needed (see our auxiliary results in the Supplementary material, and a comparative rate requirement when $\tau=0$ in e.g. Karabiyik et al.,, 2017). We present simulations under serial correlation in the Supplement, which do not show any negative effect on IC. The rest of the assumptions are standard, as they ensure weak cross-section dependence of the idiosyncratics (Assumption 2 (b)), mutual independence of factors and the error terms (Assumption 3; see also Pesaran,, 2006), and informativeness of the loadings (rank condition in Assumption 4). The loadings are deterministic, but they can also admit a random coefficient model (see e.g. De Vos and Everaert,, 2021). In our assumptions, we cover $k+1$ variables in order to accommodate MW, as well.
@@
To formulate our instrumental theoretical result, let $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\ln\mathrm{det(\overline{\mathbf{Q}}_{M})}-\ln\mathrm{det(\overline{\mathbf{Q}}_{M_{0}})}$. This quantity is important because it allows us to track the asymptotic contribution of factor estimates to the minimization problem under $M_{0}\subset M$ (over-specification) and $M\subset M_{0}$ (under-specification). Crucially, the behavior of $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ prescribes properties of $p_{N,T}$. To illustrate this, Bai, (2004) (in PC setting) shows that under over-specification, (an equivalent of) $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ is $O_{p}(1)$ when factors are non-stationary, and so the contribution of the extra estimates to the sum of squared residuals is non-negligible. This leads to $p_{N,T}\to\infty$ to heavily penalize redundant estimates. However, when factors are stationary, Bai and Ng, (2002) (and DVS) detect $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=O_{p}(C_{N,T}^{-2})$, which means that $p_{N,T}\to 0$ at a slower rate in order to penalize lightly. Proposition 1 below can be seen as the CAs equivalent of Lemma A3 and Lemma A4 of Bai, (2004).444 Strictly speaking, under-specification happens when $M\subset M_{0}$, $M_{0}\cap M\neq\emptyset$ but neither is a weak subset of each other, and when $M\cap M_{0}=\emptyset$. We only analyze the case of $M\subset M_{0}$ for brevity. As Margaritella and Westerlund, (2023) pointed out analysis would lead to the same conclusion in all the cases.
@@
  • (a)
    Under $M\subset M_{0}$, $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to_{p}+\infty$;
  • (b)
    Under $M_{0}\subset M$, $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=O_{p}(C_{N,T}^{-2})$.
Proof: Section 3 of Supplementary material.
Let us discuss Proposition 1 and start with (b). In this case, the CAs are able to approximate all $m$ factors under Assumption 4. Then $\overline{\mathbf{Q}}_{M}$ is exactly (log-determinant of) the denominator of the CCE estimator. It uses the fact that under $M_{0}\subset M$ we have $\mathbf{F}=\overline{\mathbf{X}}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}+o_{p}(1)$ for fixed $T$, and by inserting this into $\mathbf{X}_{i}$, we can show that
$\ln\mathrm{det}(\overline{\mathbf{Q}}_{M})=\ln\mathrm{det}\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}}\mathbf{V}_{i}\right)+o_{p}(1)=\ln\mathrm{det}\left(\boldsymbol{\Sigma}_{\mathbf{v}}\right)+o_{p}(1),$ (2.4)
whose argument is a positive definite matrix under general factors (see also Westerlund,, 2018). Given that $\ln\mathrm{det}(\overline{\mathbf{Q}}_{M_{0}})$ admits the same asymptotic representation, it is natural that $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ is negligible. Moreover, the rate is already detected by DVS and Bai and Ng, (2002) (for PC) under stationary $\mathbf{F}$, and it remains identical here. This rate is instructive and determines the properties of penalty $p_{N,T}$, which is the key motivation behind the functional forms in (2.2). Specifically, because redundant gmg-m CAs have an asymptotically negligible contribution to the sum of squared residuals, over-specification does not need to be penalized heavy, and so it is sufficient that $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$ as $(N,T)\to\infty$. Note that this stands in sharp contrast to findings in Lemma A4 of Bai, (2004), where excess factor estimates contribute non-trivially. Overall, the usual consistency of the CCE estimator in (1.2) under general unknown factors prevails when $M_{0}\subset M$ (see also Westerlund,, 2018, or Stauskas,, 2023).
Under $M\subset M_{0}$ in (a), the rank condition in Assumption 4 is not satisfied. This means that mgm-g factors are not asymptotically projected out, and the representation in (2.4) is invalid. Note that $\overline{\mathbf{Q}}_{M}$ is a quadratic form in $\mathbf{F}$, and we know that $\left\|\mathbf{F}^{\prime}\mathbf{F}\right\|=O_{p}(T^{1+\tau})$ as implied by Assumption 1 (a). However, the objective function offers normalization by $T$ only as the integration order is unknown. Exactly this imbalance causes divergence of $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$. A similar situation arises in Lemma A3 of Bai, (2004), where the central step is to detect the sign of divergence, and we follow this path. Note that $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\ln\det\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}\right)$, where $\boldsymbol{\Omega}_{M,M_{0}}:=T^{-\tau}[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}]\overline{\mathbf{Q}}_{M_{0}}^{-1}\geq 0$ (not zero) as $(N,T)\to\infty$, which we demonstrate in the Supplement. Using this, the difference can be simplified to

### v4

@@ Abstract
Information criteria (ICs) have been widely used in factor models to estimate an unknown number of latent factors. It has recently been shown that ICs perform well in Common Correlated Effects (CCE) and related settings when selecting a set of cross-section averages (CAs) sufficient for the factor space under stationary factors. As CAs can proxy non-stationary factors, it is tempting to claim an excellent performance of ICs under general factors, too. We show formally and in simulations that they remain consistent, but the more persistent factors are, the poorer they perform in small samples, which goes against the sentiment in the CCE/CAs literature.
@@ 1 Introduction
where $\mathbf{Z}_{i}=[\mathbf{z}_{i,1},\ldots,\mathbf{z}_{i,T}]^{\prime}\in\mathbb{R}^{T\times K}$, $\mathbf{U}_{i}$ is an error term, while $\mathbf{F}\in\mathbb{R}^{T\times m}$ are latent common factors, where $m$ denotes the number of factors, and $\mathbf{C}_{i}\in\mathbb{R}^{m\times K}$ is the loading matrix. (1.1) nests several settings. The most popular one is that of interactive effects (see Bai, 2009a, ). Then $\mathbf{Z}_{i}=[\mathbf{y}_{i},\mathbf{X}_{i}]\in\mathbb{R}^{T\times(k+1)}$, where $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ gives a model with unobserved heterogeneity. If $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}\in\mathbb{R}^{T\times k}$, we have the Common Correlated Effects (CCE) setting of Pesaran, (2006) with $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i},\boldsymbol{\gamma}_{i}$) representing idiosyncratics (loadings). (1.1) goes beyond interactive effects and CCE. Stauskas and De Vos, (2025) consider Distinct Correlated Effects (DCE), where $\mathbf{y}_{i}$ and $\mathbf{X}_{i}$ load on distinct sets of factors. Massacci and Kapetanios, (2024) use $\mathbf{Z}_{i}=\mathbf{X}_{i}$ to model a set of co-moving regressors in a forecasting equation.
@@
It is possible to estimate unobserved factors in several ways, including the Principal Components (PC) method (see Bai and Ng,, 2006) or diversified projections (see Fan and Liao,, 2022). Pesaran, (2006) suggests a simple and elegant way to proxy $\mathbf{F}$ up to a linear transformation by taking cross-section averages (CAs): $\overline{\mathbf{Z}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{Z}_{i}=\widehat{\mathbf{F}}=\mathbf{F}\overline{\mathbf{C}}+o_{p}(1)$, because CAs of the idiosyncratics are negligible under a wide range of empirically relevant assumptions (see e.g. Pesaran and Tosetti,, 2011). This is enough to estimate $\mathbf{F}$ in DCE and forecasting settings. In interactive effects, the CCE estimator of $\boldsymbol{\beta}$ is then simply the least squares (LS) estimator augmented with $\overline{\mathbf{Z}}=\widehat{\mathbf{F}}$ as an additional regressor.
@@
A substantial advantage of CAs is their ability to estimate $\mathbf{F}$ irrespective of their time series properties. Indeed, the CCE estimator enjoys popularity, because it is consistent and asymptotically normal under very general $\mathbf{F}$ in large $N,T$ settings without any modifications (see e.g. Kapetanios et al.,, 2011, Westerlund,, 2018, Westerlund and Petrova,, 2018, or Stauskas,, 2023). This property is valid if the rank of $\overline{\mathbf{C}}$ is equal to $m$, which means that the number of factors cannot exceed the effective number of CAs. However, there are reasons to ensure that the number of CAs matches $m$. In CCE, if $m<k+1$ and $TN^{-1}$ is bounded, CCE produces an asymptotic bias whose analytical correction is infeasible (see Karabiyik et al.,, 2017). This situation is common in macro datasets (see De Vos and Stauskas,, 2024). In general, too many CAs might even lead to a model that over-fits the data. In contrast, inclusion of too few generally leads to inconsistent estimates of the model parameters (see Juodis,, 2022, for a discussion). Beyond CCE, Stauskas and De Vos, (2025) show that DCE only uses $\overline{\mathbf{Z}}=\overline{\mathbf{X}}$, but $m=k$ must hold. In forecasting, the same is needed to avoid only conservative confidence intervals around factor-augmented forecasts (see Karabiyik and Westerlund,, 2021). In response to this, Margaritella and Westerlund, (2023) (MW) were the first to demonstrate that the information criteria (ICs) of Bai and Ng, (2002) and Bai, 2009b from the PC literature can be applied in the pure CCE setting under stationary factors to identify an optimal set of CAs from $\overline{\mathbf{Z}}$. De Vos and Stauskas, (2024) (DVS) show similar results when CAs are selected from $\overline{\mathbf{X}}$, which is more applicable in the DCE and forecasting settings, but the procedure has a similar theoretical basis. [^1] Even in basic CCE setting, it is advised to omit $\overline{\mathbf{y}}$ and use $\overline{\mathbf{X}}$ only to avoid computational issues (see Karavias et al.,, 2023)., [^2] PC techniques are limited in CAs setting, as they focus on $m$ by detecting the largest eigenvalues of the data matrix. They cannot detect sets of CAs as they do not have a natural ordering. With IC, we learn $m$ from the cardinality of the selected set.
@@
MW stress that an assumption of stationary $\mathbf{F}$ is required only to simplify the proofs, hinting at a much greater generality of IC. As CAs proxy a general factor structure, it is natural to evaluate this statement and re-visit the ICs proposed in both MW and DVS due to their wide applicability and similar theoretical foundation. The subtlety arises here because IC is minimized by grid-search at different combinations of CAs. Inevitably, there is a need to understand the asymptotic behavior of IC evaluated at a combination inconsistent for $\mathbf{F}$ that is non-stationary, which is a new undertaking in the CAs literature. Therefore, this study can be seen as the CAs counterpart of Bai, (2004), where classical PC results of Bai and Ng, (2002) are evaluated against pure unit root factors. Instead, we use a mildly integrated process by Magdalinos and Phillips, (2009) to model $\mathbf{F}$, which allows us to experiment with varying degrees of factor persistence. We demonstrate formally and in simulations that, while ICs remain consistent as $(N,T)\to\infty$, highly persistent factors negatively affect their small sample performance. We discuss differences between our results and those in Bai, (2004), and also explore penalty functions adapted to non-stationary $\mathbf{F}$, as suggested in the latter study. Since they make the performance of our ICs even worse, we argue that practitioners should not mechanically take recommendations from the PC literature if the goal is to select CAs. More importantly, the ICs of DVS and MW (with the original penalties) should be applied in relatively large samples if the presence of non-stationary factors is suspected.
@@ 2 Econometric Setup
We focus on the IC of DVS to analyze our theoretical results, however, the conclusions apply to MW, as well. While we provide numerical experiments on both DVS and MW for comparison, theoretical arguments regarding the latter are similar, and we relegate them to Section 3.3 of the Supplementary material. To operationalize our analysis, we introduce $M$ which is a set of column indices of $\overline{\mathbf{X}}$, and $\mathbf{q}_{M}\in\mathbb{R}^{k\times g}$ picks the corresponding $g$ averages in practice. That is, $\overline{\mathbf{X}}\mathbf{q}_{M}=\widehat{\mathbf{F}}_{M}$ defines a selection of $g$ out of $k$ CAs. Consequently, let $M_{0}$ denote the true set of averages from $\overline{\mathbf{X}}$ such that $\mathrm{rank}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}})=m$, and $|M_{0}|=m$, where $|M|$ denotes the cardinality of an arbitrary set $M$. IC under consideration is given by
@@
where $\ln(.)$ denotes the natural logarithm and $\mathrm{det}(\mathbf{A})$ is a determinant of any square matrix $\mathbf{A}$. $\overline{\mathbf{Q}}_{M}=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M}}\mathbf{X}_{i}$, $\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix, $\mathbf{A}^{+}$ is the Moore-Penrose inverse, and $p_{N,T}$ is a penalty term. [^3] In case of MW, $\overline{\mathbf{Q}}_{M}=\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M}}\widehat{\boldsymbol{\nu}}_{i}$ (scalar), where $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ and $\widehat{\boldsymbol{\beta}}$ is obtained using the CCE estimator under all available $k+1$ CAs. Then $\mathrm{IC}(M)=\ln(\overline{\mathbf{Q}}_{M})+g\cdot p_{N,T}$. Examples of feasible and most popular penalties are given by
@@
for $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$ (see more in Section 5 of Bai and Ng,, 2002). Note that (2.1) is function of (a version of) the denominator of the CCE estimator, which is robust to general unknown factors as long as CAs are (rotationally) consistent for $\mathbf{F}$ (see Theorem 1 in Westerlund,, 2018).
@@
where $|\widehat{M}|=g$ provides the estimator of the number of factors as a consequence. In DVS (and MW), it is demonstrated that $\mathbb{P}\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})<0\right)\to 0$ (therefore, $\mathbb{P}(\widehat{M}=M_{0})\to 1$) as $(N,T)\to\infty$. This result means that some other $M$ does not minimize IC asymptotically and theoretically justifies its use in the case of CAs. These findings are based on stationary $\mathbf{F}$, which has been considerably relaxed in the CCE/CAs literature. Therefore, our natural goal is to examine (2.1) in a way that departs from this restrictive assumption. Throughout our analysis, we employ the following set of assumptions.
Assumption 1. $\{\mathbf{f}_{t}\}$ is a mildly integrated process as defined in Magdalinos and Phillips, (2009), such that
$\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{G}T^{-\tau},\quad\mathbf{G}=\mathrm{diag}(g_{1},\ldots,g_{m}),\hskip 5.69054ptg_{j}\in(0,2),$
@@
Assumption 1 (a) treats the factors very flexibly and offers comparative statics, as $\tau\to 1$ increases persistence of the process. We are not interested in a specific model for $\mathbf{F}$, but Magdalinos and Phillips, (2009) allow us to vary $\tau$ and give theoretical guarantees, such as $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{F}\to_{p}\boldsymbol{\Sigma}_{\mathbf{F}}$, which is a constant positive definite matrix. The current specification ensures that we return to the usual stationarity conditions under $\tau=0$, as the eigenvalues of $\mathbf{I}_{m}-\mathbf{G}$ lie within a unit circle, so that the process can be inverted and admit a representation of $MA(\infty)$. Assumption 2 (a) is split into two parts. If the idiosyncratics $\{\mathbf{e}_{i,t}\}$ are correlated over time, for $\mathbf{E}_{i}=[\mathbf{e}_{i,1},\ldots,\mathbf{e}_{i,T}]^{\prime}$, we have that $\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}\right\|=O_{p}(T^{-\tau/2})$ (see Lemma 3.1 in Magdalinos and Phillips,, 2009), which is too slow to demonstrate that the factor estimation error is negligible. In order not to obscure the very effect of factor persistence on IC, we restrict the correlation of idiosyncratics and show that $\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}\right\|=O_{p}(T^{-(1+\tau)/2})$ as needed (see our auxiliary results in the Supplementary material, and a comparative rate requirement when $\tau=0$ in e.g. Karabiyik et al.,, 2017). We present simulations under serial correlation in the Supplement, which do not show any negative effect on IC. The rest of the assumptions are standard, as they ensure weak cross-section dependence of the idiosyncratics (Assumption 2 (b)), mutual independence of factors and the error terms (Assumption 3; see also Pesaran,, 2006), and informativeness of the loadings (rank condition in Assumption 4). The loadings are deterministic, but they can also admit a random coefficient model (see e.g. De Vos and Everaert,, 2021). In our assumptions, we cover $k+1$ variables in order to accommodate MW, as well.
Let $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\ln\mathrm{det}(\overline{\mathbf{Q}}_{M})-\ln\mathrm{det}(\overline{\mathbf{Q}}_{M_{0}})=\ln\det\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}\right)$, where $\boldsymbol{\Omega}_{M,M_{0}}:=T^{-\tau}[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}]\overline{\mathbf{Q}}_{M_{0}}^{-1}$. $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ is important since it helps describing the behavior of $\mathrm{IC}(M)$ in the neighborhood of $\mathrm{IC}(M_{0})$ under $M_{0}\subset M$ (over-specification) and $M\subset M_{0}$ (under-specification). In addition, it prescribes properties of $p_{N,T}$, which can be seen from
$\mathrm{IC}(M)-\mathrm{IC}(M_{0})=\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}+k\cdot(g-m)p_{N,T}.$ (2.4)
@@
To illustrate, Bai, (2004) (in PC setting) shows that under $M_{0}\subset M$, (an equivalent of) $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ is $O_{p}(1)$ when factors are non-stationary, and so the contribution of the extra estimates to the sum of squared residuals is non-negligible. This leads to $p_{N,T}\to\infty$ to heavily penalize redundant estimates so that $\mathrm{IC}(M)-\mathrm{IC}(M_{0})>0$. However, when factors are stationary, Bai and Ng, (2002) (and DVS) detect $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=O_{p}(C_{N,T}^{-2})$, which means that $p_{N,T}\to 0$ at a slower rate to penalize lightly to ensure $\mathrm{IC}(M)-\mathrm{IC}(M_{0})>0$ asymptotically. Proposition 1 below can be seen as the CAs equivalent of Lemma A3 and A4 of Bai, (2004). [^4] Strictly speaking, under-specification happens when $M\subset M_{0}$, $M_{0}\cap M\neq\emptyset$ but neither is a weak subset of each other, and when $M\cap M_{0}=\emptyset$. We only analyze the case of $M\subset M_{0}$ for brevity. As Margaritella and Westerlund, (2023) pointed out, analysis would lead to the same conclusion in all cases.
@@
$\textit{(a)}\hskip 5.69054pt\textit{Under}\hskip 2.84526ptM_{0}\subset M,\hskip 2.84526pt\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=O_{p}(C_{N,T}^{-2});$ (2.5)